# Jinhang Chai     Jianqing Fan

### v1

@@ 5.2 Robust Covariance Estimation
Note that there is rotational ambiguity regarding the direction of $b_{j}$ and thus $\tilde{b}_{l}$, hence an orthogonal transform is used to eliminate this ambiguity, which in our case can be taken to be a diagonal matrix with ±$\pm 1$ on the diagonal.
Remark 15.
$\lambda_{j}-\lambda_{j+1}=\Omega(p)$ is a common assumption on eigen gap (Fan et al.,, 2013, 2019). Apart from this, we don’t need the pervasiveness assumption as the magnitude of $\Sigma_{L}$ and $\Sigma_{S}$ can be arbitrary. Also, if we naively bound the estimation error of individual factor loading vector $\|b_{k}-H\widehat{b}_{k}\|$ by $\sqrt{\frac{1}{p}\sum_{j=1}^{p}\|b_{j}-H\widehat{b}_{j}\|^{2}}$, we match the same order of convergence as in Theorem 3 in Fan et al., (2019) where $v$ is of typical order $\Omega(p)$. Whether it is possible to shave a factor of $\sqrt{p}$ for bounding error of individual factor is left to future work, as our theory focuses on Frobenius norm which is an ”average-type” metric and it’s not clear whether the energy of $V-\widehat{V}$ is dispersed or not.
@@ 6 Numerical Results
In this section, we conduct extensive numerical experiments to corroborate our theoretical results. The section is divided into four experiments as follows.
@@
where $A\in\mathbb{R}^{p,r}$ and $B\in\mathbb{R}^{r,p}$ is sampled independently from $\mathcal{U}[0,1]$ in each entry. $D$ is a diagonal matrix whose diagonal entries are sampled independently from $\mathcal{U}[0,1]$. $S$ is supported on exactly $s$ entries, where the support is sampled uniformly in all $p^{2}$ entries. Conditional on the support location, entries are also drawn independently from $\mathcal{U}[0,1]$. Numerically we find that with this type of parameter generation process, the linear part is almost always of low incoherence.
The observation $Y$ is given by $Y=M+W$. We consider two types of weakly correlated noise $W$ as follows.
@@
  • Empirical probability error distribution: As the name suggests, we define the empirical probability error as the difference of the empirical probability vector and the ground-truth probability vector and its distribution as the empirical probability error distribution.
    Specifically, we consider the special case that the ground-truth probability is the uniform distribution over $\{1,2,\cdots,p\}$ so that each row of $W$ is independently drawn from the distribution of $\left(n^{-1}\sum_{i=1}^{n}\mathbf{e}_{X_{i}}-{\mathbf{1}_{p}}/{p}\right)^{\top}$, where $X_{i}$ are i.i.d. discrete random variables uniform in $\{1,2,\cdots,p\}$. [^8] As can be seen from the definition, it resembles the noise arising from the Markov transition matrix estimation problem for a special Markov transition kernel.
We apply our alternating minimization algorithm on the generated data [^9] We run the version without incoherence constraint as later we show this incoherence constraint does not matter in practice., with results shown in Figure 1. The first panel shows the case when there is no noise. The Frobenius error to the ground truth converges quickly to 0 in this case. The second panel shows the case when the noise is i.i.d. Gaussian with variance $10^{-3}$, while the third case shows the case when each row of the noise matrix is sampled independently from empirical probability error distribution. The latter two panels are based on 5 trials each. Note that trajectories vary because they are based on different realizations of the noise.
@@
no noise i.i.d. Gaussian empirical probability error
Figure 1: Trajectories of optimization errors for the alternating minimization algorithm based on synthetic data generated from 3 different distributions.
The next three experiments are conducted based on the Markov chain transition estimation problem that we consider in this paper. We generate the ground-truth transition matrix $P$ as follows [^10] The ground-truth frequency matrix $F$ can be generated accordingly based on the transition matrix $P$.,
@@
As $r$ is of minor importance, throughout the experiments we set $r=3$ as in Zhang & Wang, (2019). Also, we take the sparse level $s=p$. In the sequel, we will see that our results are nearly independent of $t$, therefore showing our method and results accommodates a wide range of scenarios with varying signal strengths of linear part and sparse part.
In implementing our proposed algorithms, we need a stopping rule. We set the maximum iteration number to be 500, and when the Frobenius distance of two sparse part matrices in two consecutive iterations differs no more than $10^{-10}$ in the Frobenius norm of the sparse matrix in the iteration, we end the loop.
@@
Figure 2: Alternating minimization for estimating Markov transition matrix. We show in the upper-left panel how error rates of estimating the frequency matrix in Frobenius norm decay with sample size. Not surprisingly, the decay rate is $\frac{1}{\sqrt{n}}$ as justified in Theorem 3.5, which is clearly shown in the upper-right panel. The lower-left panel depicts how error rates of estimating the transition matrix in $L_{1}$ norm decay in the sample size. The curve is a little wiggling, reflecting the involvement of $\pi_{\min}$ in the bound given in the Theorem 3.5. The lower right panel is shown to verify that our estimated matrix is indeed low-incoherent, in coherence with our theory.
Figure 2 shows how the error of transition matrix estimation depends on the sample size. Each point in the graph represents an average of over 20 trials. The upper-left panel depicts how error rates of estimating the frequency matrix in Frobenius norm decay with sample size. From the upper-right panel, the decay rate is approximately $n^{-1/2}$ as justified in Theorem 3.5. Also intuitively, when $t$ is larger, the elements in the transition matrix are more spiky and diverse, which results in a longer mixing time, and therefore a slightly larger error, which is reflected in the upper-right panel. The lower-left panel depicts how error rates of the transition matrix estimation in $L_{1}$ norm decay in the sample size. The curve is a little wiggling from exact linear. This is also reflected by the fact that the estimator $\widehat{P}$ is based on $\widehat{F}$ and it involves $\pi_{\min}$ in the bound given in the Theorem 3.5. The lower left panel is to verify that our estimated matrix is indeed of low incoherence, in accordance with our theory.
Experiment 3: In the third experiment, we examine the dependence of Frobenius norm error on the dimension $p$. The number of samples is fixed to be $10^{5}$ while $p$ is taken from $\{30,35,40,45,50,100,200,500\}$. We also compare two slightly different algorithmic variants as specified in Section 3.2. In particular, recall from the last section that Method2 represents the one taking into account the incoherent constraint in each iteration while Method1 is the one without the incoherent constraints, where we take $\mu=5$. Figure 3 depicts the results. They show an approximately ${1}/{\sqrt{p}}$ decay rate of the error which is coherent with our theory. Also both two panels show that including the incoherence condition in each iteration doesn’t change the performance too much.
@@
Figure 3: Alternating minimization for estimating Markov transition matrix. We plot in the left panel how error rates of estimating the frequency matrix in Frobenius norm decay with dimension $p$ using two methods. From the right panel, the decay is a typical square root $p$ as justifies in Theorem 3.5.
@@
Figure 4: The spectral etimator and the proposed Alternating Minimization. The left panel presents the Frobenius norm error of frequency matrix while the right panel depicts the $L_{1}$ norm error of the transition matrix.
@@ 7 Conclusion
In this paper, we consider the structured matrix denoising problem. In particular, the structured matrix can be written as the sum of a low-rank matrix and a sparse matrix. Our primary contribution lies in developing a deterministic error bound that does not rely on distributional assumptions, a departure from prevalent reliance on such assumptions in existing literature. The methodology we propose is an optimization procedure centered on a incoherence-constrained least-squares objective. We prove the method is optimal in a wide range of noise settings. The key intermediate step underlying the success of the constrained least-square optimization hinges on showing that the difference arbitrary two incoherent-low-rank matrices has energy spread out across entries, or cannot be too sparse. The proof of the lemma is composed of several steps starting from simple cases. Our primary motivating example revolves around the estimation of Markov chain transition kernel. The difference of the ground-truth transition matrix and empirical transition matrix, which can be regarded as the noise, is not independent across entries. Therefore, we find it necessary to relax these assumptions and consider deterministic recovery instead. The method provably attains minimax-lower bound in the class of low-rank-plus-sparse transition matrices. To solve the optimization problem, we propose an alternating minimization algorithm which alternates between optimizing the low-rank component and the sparse component. Other related problems are considered, for example, multitask regression and robust structured covariance estimation, where assumptions are relaxed compared to the factor model. Numerical results show this algorithm typically converges to the solution in a few steps, highlighting its practical utility and efficacy.
Appendix A Main Proofs
A.1 Proof of Theorem 3.5

### v2

@@ 4.2 Robust Covariance Estimation
The inequality (22) is a subspace perturbation bound. For (23), $\lambda_{j}-\lambda_{j+1}=\Omega(p)$ is a common assumption on eigen gap (Fan et al.,, 2013; Fan et al., 2019a, ). Apart from this, we don’t need the pervasiveness assumption as the magnitude of $\Sigma_{L}$ and $\Sigma_{S}$ can be arbitrary. Whether it is possible to shave a factor of $\sqrt{p}$ for the bounding error of individual factor is left to future work.
@@ 5 Numerical Results
In this section, we conduct extensive numerical simulations and real data experiments to corroborate our theoretical results.
5.1 Simulations
We first conduct four simulations as follows.
@@ 5.1 Simulations
where $A\in\mathbb{R}^{p,r}$ and $B\in\mathbb{R}^{r,p}$ are sampled independently from $\mathcal{U}[0,1]$ in each entry, $D$ is a diagonal matrix whose diagonal entries are sampled independently from $\mathcal{U}[0,1]$, $S$ is supported on exactly $s$ entries, where the support is sampled uniformly in all $p^{2}$ entries. Conditional on the support location, entries are also drawn independently from $\mathcal{U}[0,1]$. Numerically, we observe that this type of parameter generation process almost always results in the linear part with of low incoherence.
The observation $Y$ is given by $Y=M+W$. Apart from no-noise case ($W=0$), we consider two types of noise options as follows.
@@
  • Empirical distribution errors: Each row of $W$ is an independent realization of $\left(n^{-1}\sum_{i=1}^{n}\mathbf{e}_{X_{i}}-{\mathbf{1}_{p}}/{p}\right)^{\top}$, where $X_{i}$ are i.i.d. discrete random variables uniform in $[p]$.  [^9] As can be seen from the definition, it resembles the noise arising from the Markov transition matrix estimation problem for a special Markov transition kernel. Note that the entries in a row are weakly correlated, with variance in each entry given by $\frac{1}{np}(1-\frac{1}{p})$.
We apply our alternating minimization algorithm on the generated data, with different initialization methods [^10] We run the version without incoherence constraint as we later demonstrate that this constraint has minimal impact in practice., with results shown in Figure 1. The first panel shows the case when there is no noise. The Frobenius error to the ground truth converges quickly to 0 in this case. The second panel shows the case when the noise is i.i.d. Gaussian with standard deviation $10^{-3}$, while the third case shows the case when each row of the noise matrix is sampled independently from empirical probability error distribution. To facilitate fair comparison, we set $n=999$ such that the standard deviation of each entry is equal to $10^{-3}$. The latter two panels are based on 5 trials each.
We make the following observations:
  1. 1.
    All three initialization methods converge to the same optimal value. Init1 and Random Init performs similarly and are better than Init2.
  2. 2.
    In the middle panel of Figure 1, the trajectories within each initialization method appear to collapse and overlap, which aligns with the concentration property of the Gaussian ensemble.
  3. 3.
    The middle panel and the right panel are only slightly different. This is expected, as the entrywise variance is controlled to be equal for two noise settings, and the empirical probability error exhibits weak correlations across entries.
@@
no noise i.i.d. Gaussian empirical distribution error
Figure 1: Trajectories of optimization error for Algorithm 1 based on synthetic data with 3 noise scenarios. The left panel corresponds to no-noise case ($W=0$). The middle panel corresponds to the i.i.d. Gaussian noise with the standard deviation of each entry being $10^{-3}$. The right panel corresponds to the noise sampled from empirical probability error distribution with the standard error of each entry being approximately $10^{-3}$. In each panel, blue curve is the result of Init1, the red curve is the result of Init2, and the green curve is the result of Random Init. As can be seen, Init1 and Random Init performs similarly and are better than Init2. All three initialization methods converge the same optimal value.
The next three experiments are conducted based on the Markov chain transition estimation problem that we consider in this paper. For better focus, we only use the Init1 as the initialization method. We generate the ground-truth transition matrix $P$ as follows [^11] The ground-truth frequency matrix $F$ can be generated accordingly based on the transition matrix $P$.,
@@
As $r$ is of minor importance, throughout the experiments, we set $r=3$ as in Zhang & Wang, (2019). Also, we take the sparse level $s=p$. In the sequel, we will see that our results are nearly independent of $t$, therefore showing our method and results accommodate a wide range of scenarios with varying signal strengths of linear part and sparse part.
In implementing our proposed algorithms, we need a stopping rule. We set the maximum iteration number to be 3000, and when the Frobenius distance of two sparse part matrices in two consecutive iterations differs no more than $10^{-10}$ in the Frobenius norm of the sparse matrix in the iteration, we end the loop.
@@
Figure 2: Alternating minimization method for estimating Markov transition matrix. We show in the upper-left panel how error rates of estimating the frequency matrix in Frobenius norm decay with sample size. Not surprisingly, the decay rate is $\frac{1}{\sqrt{n}}$ as justified in Theorem 2.5, which is clearly shown in the upper-right panel. The lower-left panel depicts how error rates of estimating the transition matrix in $L_{1}$ norm decay in the sample size. The curve is a little wiggling, reflecting the involvement of $\pi_{\min}$ in the bound given in the Theorem 2.5. The lower right panel is shown to verify that our estimated matrix is indeed low-incoherent, consistent with our theory.
Figure 2 shows how the error of transition matrix estimation depends on the sample size. Each point in the graph represents an average of over 20 trials. The upper-left panel depicts how error rates of estimating the frequency matrix in Frobenius norm decay with sample size. From the upper-right panel, the decay rate is approximately $n^{-1/2}$ as justified in Theorem 2.5.
The lower-left panel depicts how error rates of the transition matrix estimation in $L_{1}$ norm decay in the sample size. The curve is a little wiggling from exact linear. This is also reflected by the fact that the estimator $\widehat{P}$ is based on $\widehat{F}$ and it involves $\pi_{\min}$ in the bound given in the Theorem 2.5. The lower left panel is to verify that our estimated matrix is indeed of low incoherence, in accordance with our theory.
Experiment 3: In the third experiment, we examine the dependence of Frobenius norm error on the dimension $p$. The number of samples is fixed to be $10^{5}$ while $p$ is taken from $\{30,35,40,45,50,100,200,500\}$. We also compare two slightly different algorithmic variants as specified in Section 2.2. In particular, recall from the last section that Method2 represents the one taking into account the incoherent constraint in each iteration while Method1 is the one without the incoherent constraints, where we take $\mu=5$. Figure 3 depicts the results. They show an approximately ${1}/{\sqrt{p}}$ decay rate of the error, which is coherent with our theory. Also both two panels show that including the incoherence condition in each iteration doesn’t change the performance too much.
@@
Figure 3: Alternating minimization method for estimating Markov transition matrix (continued). We plot in the left panel how error rates of estimating the frequency matrix in Frobenius norm decay with dimension $p$ using two methods. From the right panel, the decay is a typical square root $p$ as justifies in Theorem 2.5.
@@
Figure 4: The comparison of the spectral estimator and alternating minimization method in synthetic data. The left panel presents the Frobenius norm error of the frequency matrix, while the right panel depicts the $L_{1}$ norm error of the transition matrix. The estimation error of the alternating minimization method decreases monotonically as sample size increases, whereas the spectral method does not due to irreducible bias.
5.2 Real Data Experiment
We also evaluate our methodology using a real-world dataset of taxi trip records [^12] The dataset is available at https://www.nyc.gov/site/tlc/about/tlc-trip-record-data.page.. This dataset contains taxi trip information collected and provided to the NYC Taxi and Limousine Commission (TLC), including details such as pick-up and drop-off locations, trip distances, itemized fares.
For our experiment, we gathered all trips in January 2024, resulting in a total number of $2964624$ trips. For each trip, we have a one-step transition from the pick-up location to the drop-off location, both of which are districts already categorized. In preprocessing, we exclude rarely visited districts—specifically, those appearing fewer than $100$ times—leaving $74$ districts. Using this filtered data, we construct an empirical transition matrix of dimension $p=74$. Given the massive number of trips, we can assume the empirical transition matrix is sufficiently precise. Therefore, we denote that as $P^{\star}$ and treat as the ground truth for subsequent semi-synthetic experiments.
Figure 5: Singular value plot of $P^{\star}$ in real data
To continue, we examine the approximation capability of our low-rank-plus-sparse model, in comparison with low-rank model (Zhang & Wang,, 2019). We begin by plotting the singular values of $P^{\star}$ in Figure 5. The plot shows that the singular values decay to zero at around $30$-th component, suggesting that a purely low-rank assumption may not be appropriate. For our analysis, we set $r=10$ and $s=50$. For the low-rank model, we input $F^{\star}$ and compute the best approximator $F_{L}$, which is equivalent to the top $r$ singular components of $F^{\star}$. We can then conduct projection and row-wise normalization to derive $P_{L}$. For the low-rank-sparse model, we input $F^{\star}$ to Algorithm 1 to obtain the best approximator $F_{LS}$, from which we derive $P_{LS}$ in a similar way. The approximation results are shown in Table 1. Compared to the low-rank model, the low-rank-plus-sparse model achieves a significantly better fit to the real data, substantially reducing the approximation error.
$\|F^{\star}-F_{LS}\|_{F}$ 0.0037 $\|F^{\star}-F_{L}\|_{F}$ 0.0349
$\|P^{\star}-P_{LS}\|_{1}/p$ 0.0075 $\|P^{\star}-P_{L}\|_{1}/p$ 0.0363
Table 1: Approximation results in real data (subscripts $L$ and $LS$ represent low-rank, and low-rank-plus-sparse model.)
Furthermore, we conduct semi-synthetic experiments, generating Markov chain data regarding $P^{\star}$ as the ground truth transition matrix. For sample size $n$ in $\{5000k:1\leq k\leq 20\}$, we generate a single trajectory of length $n$. We compare our alternating minimization method to the spectral method, with results averaged over $100$ trials. The outcomes are plotted in Figure 6. Similar to Figure 4, our alternating minimization method significantly outperforms the spectral method. Notably, the estimation error of the spectral method does not decrease as the sample size increases, indicating the presence of irreducible bias. In contrast, the low-rank-plus-sparse model, along with our method, aligns more closely with the data, achieving better overall performance.
Refer to caption Refer to caption
Figure 6: The Comparison of spectral estimator (in blue) and alternating minimization method (in black) in real data. The left panel presents the Frobenius norm error of the frequency matrix, while the right panel depicts the $L_{1}$ norm error of the transition matrix. The estimation error of the alternating minimization method decreases approximately monotonically as sample size increases, whereas the spectral method does not due to irreducible bias.
@@ 6 Conclusion
In this paper, we consider the structured matrix denoising problem. In particular, the structured matrix can be written as the sum of a low-rank matrix and a sparse matrix. Our primary contribution lies in developing a deterministic error bound that does not rely on distributional assumptions, a departure from prevalent reliance on such assumptions in existing literature. The methodology we propose is an optimization procedure centered on an incoherence-constrained least-squares objective. We prove the method is optimal in a wide range of noise settings. The key intermediate step underlying the success of the constrained least-square optimization hinges on showing that the difference between two arbitrary incoherent-low-rank matrices has energy spread out across entries or cannot be too sparse. The proof of the lemma is composed of several steps, starting from simple cases. Our primary motivating example revolves around the estimation of the Markov chain transition kernel. The difference of the ground-truth transition matrix and empirical transition matrix, which can be regarded as the noise, is not independent across entries. Therefore, we find it necessary to relax these assumptions and consider deterministic recovery instead. The method provably attains the minimax-lower bound in the class of low-rank-plus-sparse transition matrices. To solve the optimization problem, we propose an alternating minimization algorithm that alternates between optimizing the low-rank component and the sparse component. Other related problems are considered, for example, multitask regression and robust structured covariance estimation, where assumptions are relaxed compared to the factor model. Numerical results show this algorithm typically converges to the solution in a few steps, highlighting its practical utility and efficacy.
Supplementary Material
In this supplementary material, Appendix A provides a discussion of the related works. Appendix B contains the main proofs of results in Section 2, while Appendix C includes the proofs related to Section 3. We then elaborate on the proof of the key separation lemma in Appendix D. We include in Appendix E the proofs in Section 4. Appendix F is devoted to the extension of rectangular estimand. Several technical lemmas are listed in the Appendix G.
Appendix A Related work
Our work is related to the following lines of research directions.
  • Noisy Matrix Recovery: Noisy matrix recovery has been studied intensively in different settings such as noisy matrix completion (Candes & Plan,, 2010; Candes & Recht,, 2012; Gross,, 2011; Chen et al.,, 2019; Chen et al., 2020b, ) and robust PCA (Candès et al.,, 2011; Chandrasekaran et al.,, 2011; Chen et al., 2021b, ). There are mainly two popular ways to formulate the estimation problem, namely convex relaxation (Negahban & Wainwright,, 2012; Koltchinskii et al.,, 2011; Chen et al., 2020b, ; Chen et al., 2021b, ) and nonconvex Burer–Monteiro approach with gradient descent (Sun & Luo,, 2016; Ma et al.,, 2018; Chen et al., 2020a, ) or projected descent (Chen & Wainwright,, 2015; Zheng & Lafferty,, 2016). See overview papers (Chi et al.,, 2019; Chen & Chi,, 2018). With no missing entries, low-rank-and-sparse noisy matrix recovery has been studied in Hsu et al., (2011); Agarwal et al., (2012); Chandrasekaran et al., (2011); Netrapalli et al., (2014). Among those, one of the most relevant works to us is Agarwal et al., (2012), where the problem of study is under the name of noisy matrix decomposition. Other related matrix or vector sensing problems include blind deconvolution (Ma et al.,, 2018; Chen et al.,, 2023), phase retrieval (Ma et al.,, 2018), tensor decomposition (Ge & Ma,, 2017), etc.
  • Markov Transition Estimation: There has been a flurry of research on transition estimation. Some works consider the plain tabular case with no statistical structures imposed, such as Hao et al., (2018), Wolfer & Kontorovich, (2019), and Wolfer & Kontorovich, (2021). However, the rate inevitably depends on the size of the state space, which can be huge. To mitigate this dimension issue, Zhang & Wang, (2019) imposes low-rankness and derives the minimax rate of estimation; see also Zhu et al., (2022). Section 3 in Wolfer, (2024) considers a sparse transition matrix and presents improved estimation results. To the best of our knowledge, our work is the first to consider the low-rank-plus-sparse structure on the transition matrix.
  • Spectral methods: Spectral methods have become a paradigm for many statistical machine problems (Chen et al., 2021a, ), and have been a go-to choice for various kinds of problems, including PCA and factor analysis (Johnstone & Paul,, 2018; Fan et al.,, 2018), community detection (Abbe,, 2017), clustering (Ndaoud,, 2018; Löffler et al.,, 2021), ranking (Chen & Suh,, 2015), covariance estimation (Fan et al.,, 2013), etc. Although our approach does not utilize eigenvector perturbation theory due to the existence of a potentially large sparse part, the techniques bear a lot of similarity with the spectral methods, which intensively use singular value decomposition and matrix analysis.
Here we give a detailed comparison of a few more related literature. Agarwal et al., (2012) bears the most resemblance to our work, which studied in a similar setting. Notably, their intermediate deterministic bound also accounts for arbitrary noise. However, they imposed a spikiness condition that each entry of the low-rank part of the ground-truth matrix is upper bounded in magnitude and the final deterministic bound involves a trailing term as appeared in Table 2. A huge drawback is that this bound results in inconsistency where the estimation error is lower bounded from zero even with infinite samples. Our result improves this.
Hsu et al., (2011) also considered the arbitrary noisy setting and gave bounds on estimating errors. However, compared to our work, there are several differences. Firstly, their bound is not tight in general. In particular, their bound on low-rank part is not optimal in their Theorem 3 and they can’t get optimal Frobenius norm bound while ours can. Secondly, our results apply to every linear observation model with restricted strong convexity while their results only apply to the standard low-rank-plus-sparse matrix estimation. Thirdly, unlike their work, we don’t need prior knowledge about the error scale. We have less hyperparameters to tune than theirs. Again, our conditions on the sparse part are more relaxed than theirs.
There is a more recent work on Robust PCA (Chen et al., 2021b, ), which considered the setting where noise is i.i.d. and sub-Gaussian across entries. As previously mentioned, our results don’t need such an assumption. Their techniques of analyzing the rate convergence of gradient descent using leave-one-out highly rely on the assumption of independence of noise across entries. Moreover, unlike theirs, we can discard the assumption on the randomness of the sparse locations, which is considered fixed in our setting.
In contrast to the aforementioned papers that considered noisy observations, Chandrasekaran et al., (2011) and Candès et al., (2011) studied under the noiseless setting where they proved with high probability that exact recovery is possible. In contrast, our setting involves noise, and therefore, fully recovering the true signal is impossible. Our conditions on the sparse part are more stringent than theirs.
Note that the key separation lemma D.1 characterizes the sparsity of arbitrary difference of two incoherent low-rank matrices. One can impose a more fine-grained structure on sparsity such as row-wise or column-wise sparsity (Chandrasekaran et al.,, 2011) and prove similar separation results. This is also an interesting future direction.
Appendix B Main Proofs in Section 2
B.1 Proof of Theorem 2.5